# A possible connection between quantum mechanics and spacetime

### v3

@@
Hong Wanga, Jin Wangb,111jin.wang.1@stonybrook.edu aState Key Laboratory of Electroanalytical Chemistry, Changchun Institute of Applied Chemistry, Chinese Academy of Sciences, Changchun 130022, China
bDepartment of Chemistry and Department of Physics and Astronomy, State University of New York at Stony Brook, NY 11794, USA
Recent developments in holographic gravity suggest that spacetime structure may be deeply related to quantum mechanics. In this work, from a different perspective, we demonstrate that the wave-particle duality can be interpreted as the uncertainty of spacetime for the particle. Summarizing all possible trajectories in conventional path integral quantum mechanics can be transformed into the summation of all possible spacetime metrics. Furthermore, we emphasize that in conventional quantum gravity, it is possible that the classical matter fields correspond to the quantum spacetime. We argue that this is not quite reasonable and propose a new path integral quantum gravity model based on the new interpretation of wave-particle duality. In this model, the aforementioned drawback of conventional quantum gravity naturally disappears.
@@ 1 Introduction
There are two equivalent methods for quantizing a system: the path integral and canonical quantization. In path integral quantum mechanics [23], to calculate the transition amplitude, one must sum up all possible trajectories. The trajectory with the minimum action is the classical trajectory. For a free particle, it is a straight line. Other trajectories are not straight lines and do not satisfy the Hamilton-Jacobi equation. Only the classical trajectory corresponds clearly to a dynamical equation (Hamilton-Jacobi equation). Thus, the classical trajectory possesses a special status. However, it may be more natural to consider all trajectories as solutions of some dynamical equations. What is this relevant dynamical equation? Is this idea feasible, and what implications does it have for quantum mechanics?
In this work, we show that different trajectories in the path integral can be regarded as solution of the Hamilton-Jacobi equation in different spacetimes. For the free particle, different trajectories in the path integral can be interpreted as the geodesics in different spacetimes. We find that the integration over all possible trajectories in the path integral quantum mechanics can be transformed into the integration over various spacetime metrics. This results in the spacetime representation of quantum mechanics. We show that in the spacetime representation, when using the path integral to quantize any system, one only needs to set the spacetime metric as the integral variable. We demonstrate that the wave-particle duality can be interpreted as the uncertainty of the spacetime structure for the microscopic particle. Thus, the quantum characteristics of any system can be attributed to the uncertainty of the spacetime structure. This result suggests that quantum mechanics is deeply connected to spacetime structure. Recent developments in holographic gravity also suggest a deep relationship between quantum mechanics and spacetime [24, 25, 26, 27, 28].
In conventional path integral quantum gravity, classical matter fields may correspond to the quantum spacetime [29]. We demonstrate that this is not quite reasonable. Based on the new interpretation of the wave-particle duality, we propose that in path integral quantum gravity, only the spacetime metric should be considered as the integral variable. By adopting this proposal, the aforementioned problem in quantum gravity can be easily circumvented. Throughout this study, we work in units where $G=\hbar=c=1$.
@@ 2.1 A single particle
In quantum mechanics, there are still conceptual puzzles regarding the wave-particle duality and the wave function. The evolution of the wave function is governed by the Schrödinger equation, which can be expressed in various representations. Different representations are connected through a unitary transformation. In this section, we will explore the possibility of a spacetime representation in quantum mechanics, which may help to understand the wave-particle duality and the wave function.
The spacetime coordinates can be chosen arbitrarily. Cartesian coordinates are convenient for our illustration. We will use symbols $X^{\mu}$, $X^{\alpha}$, $X^{\beta}$, … to represent Cartesian coordinates. Other coordinates will be denoted by different symbols, such as $x^{\mu}$, $(\theta,\varphi)$, …. In Cartesian coordinates, the spacetime metric is
$ds^{2}=g_{\alpha\beta}dX^{\alpha}dX^{\beta}.$ (1)
Various spacetimes correspond to different $g_{\alpha\beta}$. When the spacetime is flat, Cartesian coordinates imply that $g_{\alpha\beta}=\eta_{\alpha\beta}$, where $\eta_{\alpha\beta}=diag(1,-1,-1,-1)$ represents the Minkowski metric.
@@
where, $m_{0}$ and $g_{\alpha\beta}$ represent the rest mass of the particle and the spacetime metric, respectively. $\tau$ is the proper time parameter. Taking $\delta S/\delta X^{\mu}=0$, one can obtain the classical dynamical equation (geodesic equation) of the particle as
@@
Here, $\Gamma^{\mu}_{\alpha\beta}$ is the connection. For the torsionless spacetime, $\Gamma^{\mu}_{\alpha\beta}=\frac{1}{2}g^{\mu\rho}(g_{\rho\alpha,\beta}+g_{\rho\beta,\alpha}-g_{\alpha\beta,\rho})$.
In a specific spacetime $g_{\alpha\beta}$, if the initial state of the particle is fixed, then equation (3) has a unique solution. Changing the metric $g_{\alpha\beta}$ continuously, the solution may also be changed continuously. Thus, the trajectory of the particle can be seen as a functional of the metric. We formally denote the solution of the geodesic equation (3) as
@@
Equation (3) is a nonlinear equation. It is difficult to give a general solution for the geodesic equation. Thus, the explicitly form of the functional $\mathcal{F}^{\mu}$ is difficult to specify. However, the conclusions of this work are independent of the detailed form of the functional $\mathcal{F}^{\mu}$. We point out that the metric $g_{\alpha\beta}$ and the solution of equation (3) is not in a one-to-one correspondence. Fixing the metric $g_{\alpha\beta}$ and the initial state of the particle, there is only one trajectory that satisfies equation (3). Conversely, given the trajectory and the initial state of the particle, there are infinite spacetime metrics are compatible with the trajectory of the particle. In other words, one can not determine the spacetime structure (metric) according to a trajectory of the free particle.
@@
Equation (3) can also be interpreted as a particle moving in a gravitational field. When the gravitational field strength is zero, the trajectory of the particle is a straight line. By changing the gravitational field, the trajectory can take on any curve. Thus, it is reasonable to infer that for any given curve (not violating physical principles), there exists a gravitational field that makes it to become the particle’s trajectory. In other words, given any one physically allowed curve, it can always be embedded into a spacetime where it becomes a geodesic. By changing the local spacetime structure far from the trajectory of the particle, the dynamics of the particle remain unaffected. Hence, for any physically allowed curve, there are infinite types of spacetime where one of the geodesics matches this curve.
@@
In the Minkowski spacetime, according to conventional path integral quantum mechanics, the transition amplitude of a relativistic particle from the initial state $|X^{\mu}(\tau_{i})\rangle$ to the final state $|X^{\mu}(\tau_{f})\rangle$ can be written as
$\langle X^{\mu}(\tau_{f})|X^{\mu}(\tau_{i})\rangle=\mathscr{N}\int DX^{\mu}(\tau)\ \mathrm{exp}\{im_{0}\int_{\tau_{i}}^{\tau_{f}}d\tau\sqrt{\eta_{\alpha\beta}\frac{dX^{\alpha}}{d\tau}\frac{dX^{\beta}}{d\tau}}\}.$ (5)
@@
Here, $\mathscr{N}$ represents the normalized constant. $\int DX^{\mu}(\tau)$ represents the summation over all trajectories connecting the initial and final states. The factor $\int_{\tau_{i}}^{\tau_{f}}d\tau\sqrt{\eta_{\alpha\beta}\frac{dX^{\alpha}}{d\tau}\frac{dX^{\beta}}{d\tau}}$ represents the length of the trajectory $X^{\mu}(\tau)$ between the initial and final state. The lengths of different trajectories may be different. The classical trajectory is the solution to the Hamilton-Jacobi equation, while all other trajectories do not satisfy the classical dynamical equation. According to the conventional interpretation, the spacetime background in the path integral (5) is deterministic (Minkowski spacetime). Various trajectories in equation (5) are curves within the Minkowski spacetime.
@@
Every trajectory in the path integral (5) is a possible path of the particle from the initial state $|X^{\mu}(\tau_{i})\rangle$ to the final state $|X^{\mu}(\tau_{f})\rangle$. And each trajectory should be physically allowed. As discussed previously, for any given trajectory $X^{\mu}(\tau)$ in the path integral (5), there exists a spacetime where this trajectory is a geodesic. (This point is crucial for this study. We provide a more rigorous demonstration of this point in the appendix.) Therefore, the length of the trajectory can be written as
@@
A simple example is that when calculating the length of a circle defined by $(X^{1})^{2}+(X^{2})^{2}=1$, one can view the circle as an ordinary curve in the Euclidean space $ds^{2}=(dX^{1})^{2}+(dX^{2})^{2}$, and use the formula $2\int_{-1}^{1}\sqrt{1+(dX^{2}/dX^{1})^{2}}dX^{1}$ to calculate its length. One can also regard the circle as a geodesic on the two dimensional sphere surface space $ds^{2}=g_{ij}dx^{i}dx^{j}=d\theta^{2}+\mathrm{sin}^{2}\theta d\varphi^{2}$, and then use the formula $\int_{\tau_{i}}^{\tau_{f}}d\tau\sqrt{g_{ij}\frac{dx^{i}}{d\tau}\frac{dx^{j}}{d\tau}}$ to calculate its length. Thus, one can obtain the equation
@@
The local spacetime structure far away from the circle has no influence on its properties. Thus, one can find infinity types of $g_{ij}$ that satisfy equation (7).
For Minkowski spacetime, only the straight line between the initial state $|X^{\mu}(\tau_{i})\rangle$ and the final state $|X^{\mu}(\tau_{f})\rangle$ is the geodesic. In a curved spacetime, it is possible that more than one trajectories in the path integral (5) can be interpreted as the geodesics in the same spacetime. However, in this work, we consistently treat different trajectories in the path integral (5) as geodesics in distinct spacetimes. Each trajectory corresponds to a different spacetime. Thus, the integral over different paths in equation (5) can be transformed into an integral over different metrics. Therefore, using equations (4) and (6), equation (5) can be written as
$\begin{split}\langle X^{\mu}(\tau_{f})|X^{\mu}(\tau_{i})\rangle&=\mathscr{N}\int DX^{\mu}\ \mathrm{exp}\{im_{0}\int_{\tau_{i}}^{\tau_{f}}d\tau\sqrt{\eta_{\alpha\beta}\frac{dX^{\alpha}}{d\tau}\frac{dX^{\beta}}{d\tau}}\}\\ &=\mathscr{N}\int D\mathcal{F}^{\mu}\ \mathrm{exp}\{im_{0}\int_{\tau_{i}}^{\tau_{f}}d\tau\sqrt{g_{\alpha\beta}\frac{d\mathcal{F}^{\alpha}}{d\tau}\frac{d\mathcal{F}^{\beta}}{d\tau}}\}\\ &=\mathscr{N}\int Dg_{\mu\nu}\ \mathcal{J}\ \mathrm{exp}\{im_{0}\int_{\tau_{i}}^{\tau_{f}}d\tau\sqrt{g_{\alpha\beta}\frac{d\mathcal{F}^{\alpha}}{d\tau}\frac{d\mathcal{F}^{\beta}}{d\tau}}\}.\end{split}$ (8)
@@
Noted that for every trajectory $X^{\mu}(\tau)$, there are infinitely many types of metric $g_{\alpha\beta}$ that satisfy equation (6). We define these spacetimes as the same class of spacetime and use $g_{\alpha\beta}]$ to denote it. Every trajectory in the path integral (5) corresponds to different classes of spacetime $g_{\alpha\beta}]$. For a specific spacetime, it belongs to and only belongs to one of the classes $g_{\alpha\beta}]$. As each trajectory is only counted once in the path integral (5), thus, in each class of spacetime $g_{\alpha\beta}]$, only one metric contributes to the path integral. However, in equation (8), the integral $\int Dg_{\mu\nu}$ contains all metrics for every class of spacetime $g_{\alpha\beta}]$. We need to eliminate the redundant metrics.
@@
$\mathcal{G}^{(m)}(g_{\alpha\beta})=0,\ \ \ \ \ \ \ \ \ m=1,\ 2,\ ...,\ M.$ (9)
@@
Here, $M$ represents the total number of gauge conditions. Similar to the scenario with Yang-Mills gauge fields, the gauge conditions have a certain degree of arbitrariness. As long as the solution to equations (9) can uniquely determine a metric from each class of spacetime $g_{\alpha\beta}]$, the gauge conditions are considered valid. For every trajectory in the path integral (5), there exists a single metric that satisfies both equations (6) and (9).
@@
According to the Faddeev-Popov method [31, 32], to eliminate the redundant metrics, equation (8) should be modified as
@@
$\langle X^{\mu}(\tau_{f})|X^{\mu}(\tau_{i})\rangle=\mathscr{N}\int Dg_{\mu\nu}\ \mathcal{J}\ \Delta(\mathcal{G}^{(m)})\prod_{m=1}^{M}\delta(\mathcal{G}^{(m)})\ \mathrm{exp}\{im_{0}\int_{\tau_{i}}^{\tau_{f}}d\tau\sqrt{g_{\alpha\beta}\frac{d\mathcal{F}^{\alpha}}{d\tau}\frac{d\mathcal{F}^{\beta}}{d\tau}}\}.$ (10)
@@
Here, $\Delta(\mathcal{G}^{(m)})$ and $\delta(x)$ represent the Faddeev-Popov determinant and the $\delta$-function, respectively. The Jacobian determinant $\mathcal{J}$ is determined by equation (4), while the Faddeev-Popov determinant depends on the gauge conditions (9). The conclusions of this work are independent of the explicit form of $\mathcal{J}$, $\Delta(\mathcal{G}^{(m)})$ and $\mathcal{F}^{\mu}$. Thus, we will avoid solving them. This is a challenging issue that needs careful study in the future.
According to the conventional interpretation of quantum mechanics, the path integral in equation (5) includes some discrete or non-differentiable trajectories. One may argue that such trajectories cannot be regarded as geodesics in a smooth and continuous spacetime. However, we believe that our demonstration in the appendix can be extended to non-smooth or non-continuous spacetimes. This suggests that the discrete or non-differentiable trajectories in equation (5) can still be interpreted as geodesics in a certain class of non-smooth or non-continuous spacetime. If these trajectories are physically allowed, then the path integral in equation (10) should involve contributions from such non-smooth or non-continuous spacetimes. Nevertheless, to our knowledge, the existence of these trajectories has yet to be directly verified experimentally. Furthermore, the mathematics of non-smooth or non-continuous spacetimes is complicated. For these reasons, we will not explore these trajectories in detail in the present work.
@@
Equation (10) represents a new kind of representation of quantum mechanics. We name it the spacetime representation. Equation (5) is equivalent to equation (10). Thus, the integration over different trajectories can be transformed into integration over various spacetime metrics. Equation (10) indicates that the quantum characteristics of the particle can be attributed to the uncertainty of the spacetime structure for the microscopic particle. In the conventional interpretation of path integral quantum mechanics, the spacetime structure is deterministic, and only the classical trajectory corresponds to the Hamilton-Jacobi equation. All other trajectories are not solutions of any dynamical equations. Thus, the classical trajectory possesses a special status. However, we argued that for the free particle, all trajectories can be seen as the geodesics in different spacetimes. Thus, different trajectories are solutions of the geodesic equation in different spacetimes. This perspective offers a more natural understanding of these trajectories. In table 1, we outlined the key distinctions between the conventional path integral and the spacetime representation for a single relativistic free particle. When $m_{0}\rightarrow\infty$, the microscopic particle becomes a macroscopic object. The coherence disappears and the dynamics of the particle can be described by the classical physics. Consequently, the spacetime structure becomes deterministic for macroscopic object.
  Conventional path integral Spacetime representation
Partition
function
Equation (5) Equation (10)
Integral
variable
Particle’s trajectory Spacetime metric
Spacetime
structure
Spacetime structure is deterministic (Minkowski spacetime). Spacetime structure is uncertain for the microscopic particle.
Particle’s
trajectories
@@
Only the classical trajectory is the solution of the Hamilton-Jacobi equation. All other trajectories in the path integral (5) do not correspond to any dynamical equation. All trajectories in the path integral (5) are viewed as geodesics in different spacetimes.

### v4

@@
Hong Wanga, Jin Wangb,c,111jin.wang.1@stonybrook.edu aState Key Laboratory of Electroanalytical Chemistry, Changchun Institute of Applied Chemistry, Chinese Academy of Sciences, Changchun 130022, China
bCenter of Theoretical Physics, College of Physics, Jilin University, Changchun 130012, China
cDepartment of Chemistry and Department of Physics and Astronomy, State University of New York at Stony Brook, NY 11794, USA
Recent developments in holographic gravity suggest that spacetime structure may be deeply related to quantum mechanics. In this work, from a different perspective, we demonstrate that wave–particle duality can be interpreted as the uncertainty of spacetime for the particle. Summarizing all possible trajectories in conventional path integral quantum mechanics can be transformed into the summation of all possible spacetime metrics. Furthermore, we emphasize that in conventional quantum gravity, it is possible that the classical matter fields correspond to quantum spacetime. We argue that this is not quite reasonable and propose a new path integral quantum gravity model based on the new interpretation of wave–particle duality. In this model, the aforementioned drawback of conventional quantum gravity naturally disappears.
@@ 1 Introduction
There are two equivalent methods for quantizing a system: the path integral and canonical quantization. In path integral quantum mechanics [23], to calculate the transition amplitude, one must sum up all possible trajectories. The trajectory with the minimum action is the classical trajectory. For a free particle, it is a straight line. Other trajectories are not straight lines and do not satisfy the Hamilton–Jacobi equation. Only the classical trajectory corresponds clearly to a dynamical Equation (Hamilton–Jacobi equation). Thus, the classical trajectory possesses a special status. However, it may be more natural to consider all trajectories solutions of some dynamical equations. What is this relevant dynamical equation? Is this idea feasible, and what implications does it have for quantum mechanics?
In this work, we show that different trajectories in the path integral can be regarded as solutions of the Hamilton–Jacobi equation in different spacetimes. For the free particle, different trajectories in the path integral can be interpreted as the geodesics in different spacetimes. We find that the integration over all possible trajectories in the path integral quantum mechanics can be transformed into the integration over various spacetime metrics. This results in the spacetime representation of quantum mechanics. We show that in the spacetime representation, when using the path integral to quantize any system, one only needs to set the spacetime metric as the integral variable. We demonstrate that wave–particle duality can be interpreted as the uncertainty of the spacetime structure for the microscopic particle. Thus, the quantum characteristics of any system can be attributed to the uncertainty of the spacetime structure. This result suggests that quantum mechanics is deeply connected to spacetime structure. Recent developments in holographic gravity also suggest a deep relationship between quantum mechanics and spacetime [24, 25, 26, 27, 28].
In conventional path integral quantum gravity, classical matter fields may correspond to the quantum spacetime [29]. We demonstrate that this is not quite reasonable. Based on the new interpretation of wave–particle duality, we propose that in path integral quantum gravity, only the spacetime metric should be considered the integral variable. By adopting this proposal, the aforementioned problem in quantum gravity can be easily circumvented. Throughout this study, we work in units where $G=\hbar=c=1$.
@@ 2.1 A Single Particle
In quantum mechanics, there are still conceptual puzzles regarding wave–particle duality and the wave function. The evolution of the wave function is governed by the Schrödinger equation, which can be expressed in various representations. Different representations are connected through a unitary transformation. In this section, we will explore the possibility of a spacetime representation in quantum mechanics, which may help to understand wave–particle duality and the wave function.
The spacetime coordinates can be chosen arbitrarily. Cartesian coordinates are convenient for our illustration. We will use symbols $X^{\mu}$, $X^{\alpha}$, $X^{\beta}$, …to represent Cartesian coordinates. Other coordinates will be denoted by different symbols, such as $x^{\mu}$, $(\theta,\varphi)$, …. We denote the spacetime metric as $g_{\alpha\beta}$. When the spacetime is flat, Cartesian coordinates imply that $g_{\alpha\beta}=\eta_{\alpha\beta}$, where $\eta_{\alpha\beta}=diag(1,-1,-1,-1)$ represents the Minkowski metric.
@@
where $m_{0}$ represents the rest mass of the particle. The negative sign is explicitly factored out to facilitate the discussion of dynamics in the non-relativistic limit. $\tau$ is the proper time parameter. Taking $\delta S/\delta X^{\mu}=0$, one can obtain the classical dynamical Equation (geodesic equation) of the particle as
@@
Here, $\Gamma^{\mu}_{\alpha\beta}$ is the connection.
In a specific spacetime $g_{\alpha\beta}$, if the initial state of the particle is fixed, then Equation (2) has a unique solution. Changing the metric $g_{\alpha\beta}$ continuously, the solution may also be changed continuously. Thus, the trajectory of the particle can be seen as a functional of the metric. We formally denote the solution to the geodesic Equation (2) as
@@
Equation (2) is a nonlinear equation. It is difficult to give a general solution for the geodesic equation. Thus, the explicit form of the functional $\mathcal{F}^{\mu}$ is difficult to specify. However, the conclusions of this work are independent of the detailed form of the functional $\mathcal{F}^{\mu}$. We point out that the metric $g_{\alpha\beta}$ and the solution to Equation (2) are not in a one-to-one correspondence. Fixing the metric $g_{\alpha\beta}$ and the initial state of the particle, there is only one trajectory that satisfies Equation (2). Conversely, given the trajectory and the initial state of the particle, there are infinite spacetime metrics that are compatible with the trajectory of the particle. In other words, one can not determine the spacetime structure (metric) according to the trajectory of the free particle.
@@
Equation (2) can also be interpreted as a particle moving in a gravitational field. When the gravitational field strength is zero, the trajectory of the particle is a straight line. By changing the gravitational field, the trajectory can take on any curve. Thus, it is reasonable to infer that for any given curve (not violating physical principles), there exists a gravitational field that causes it to become the particle’s trajectory. In other words, given any one physically allowed curve, it can always be embedded into a spacetime where it becomes a geodesic. By changing the local spacetime structure far from the trajectory of the particle, the dynamics of the particle remain unaffected. Hence, for any physically allowed curve, there are infinite types of spacetime where one of the geodesics matches this curve.
@@
In Minkowski spacetime, according to conventional path integral quantum mechanics, the transition amplitude of a relativistic particle from the initial state $|X^{\mu}(\tau_{i})\rangle$ to the final state $|X^{\mu}(\tau_{f})\rangle$ can be written as
$\langle X^{\mu}(\tau_{f})|X^{\mu}(\tau_{i})\rangle=N\int DX^{\mu}(\tau)\ \mathrm{exp}\{im_{0}\int_{\tau_{i}}^{\tau_{f}}d\tau\sqrt{\eta_{\alpha\beta}\frac{dX^{\alpha}}{d\tau}\frac{dX^{\beta}}{d\tau}}\}.$ (4)
@@
Here, $N$ represents the normalized constant. $\int DX^{\mu}(\tau)$ represents the summation over all trajectories connecting the initial and final states. The factor $\int_{\tau_{i}}^{\tau_{f}}d\tau\sqrt{\eta_{\alpha\beta}\frac{dX^{\alpha}}{d\tau}\frac{dX^{\beta}}{d\tau}}$ represents the length of the trajectory $X^{\mu}(\tau)$ between the initial and final state. The lengths of different trajectories may be different. The classical trajectory is the solution to the Hamilton–Jacobi equation, while all other trajectories do not satisfy the classical dynamical equation. According to the conventional interpretation, the spacetime background in the path integral (4) is deterministic (Minkowski spacetime). Various trajectories in Equation (4) are curves within Minkowski spacetime.
@@
Every trajectory in the path integral (4) is a possible path of the particle from the initial state $|X^{\mu}(\tau_{i})\rangle$ to the final state $|X^{\mu}(\tau_{f})\rangle$. Furthermore, each trajectory should be physically allowed. As discussed previously, for any given trajectory $X^{\mu}(\tau)$ in the path integral (4), there exists a spacetime where this trajectory is a geodesic. (This point is crucial for this study. We provide a more rigorous demonstration of this point in Appendix A.) Therefore, the length of the trajectory can be written as
@@
A simple example is that when calculating the length of a circle defined by $(X^{1})^{2}+(X^{2})^{2}=1$, one can view the circle as an ordinary curve in the Euclidean space $ds^{2}=(dX^{1})^{2}+(dX^{2})^{2}$, and use the formula $2\int_{-1}^{1}\sqrt{1+(dX^{2}/dX^{1})^{2}}dX^{1}$ to calculate its length. One can also regard the circle as a geodesic on the two-dimensional sphere surface space $ds^{2}=g_{ij}dx^{i}dx^{j}=d\theta^{2}+\mathrm{sin}^{2}\theta d\varphi^{2}$, and then use the formula $\int_{\tau_{i}}^{\tau_{f}}d\tau\sqrt{g_{ij}\frac{dx^{i}}{d\tau}\frac{dx^{j}}{d\tau}}$ to calculate its length. Thus, one can obtain the equation
@@
The local spacetime structure far away from the circle has no influence on its properties. Thus, one can find infinite types of $g_{ij}$ that satisfy Equation (6).
For Minkowski spacetime, only the straight line between the initial state $|X^{\mu}(\tau_{i})\rangle$ and the final state $|X^{\mu}(\tau_{f})\rangle$ is the geodesic. In curved spacetime, it is possible that more than one trajectory in the path integral (4) can be interpreted as the geodesics in the same spacetime. However, in this work, we consistently treat different trajectories in the path integral (4) as geodesics in distinct spacetimes. Each trajectory corresponds to a different spacetime. Thus, the integral over different paths in Equation (4) can be transformed into an integral over different metrics. Therefore, using Equations (3) and (5), Equation (4) can be written as
$\begin{split}\langle X^{\mu}(\tau_{f})|X^{\mu}(\tau_{i})\rangle&=N\int DX^{\mu}\ \mathrm{exp}\{im_{0}\int_{\tau_{i}}^{\tau_{f}}d\tau\sqrt{\eta_{\alpha\beta}\frac{dX^{\alpha}}{d\tau}\frac{dX^{\beta}}{d\tau}}\}\\ &=N\int D\mathcal{F}^{\mu}\ \mathrm{exp}\{im_{0}\int_{\tau_{i}}^{\tau_{f}}d\tau\sqrt{g_{\alpha\beta}\frac{d\mathcal{F}^{\alpha}}{d\tau}\frac{d\mathcal{F}^{\beta}}{d\tau}}\}\\ &=N\int Dg_{\mu\nu}\ \mathcal{J}\ \mathrm{exp}\{im_{0}\int_{\tau_{i}}^{\tau_{f}}d\tau\sqrt{g_{\alpha\beta}\frac{d\mathcal{F}^{\alpha}}{d\tau}\frac{d\mathcal{F}^{\beta}}{d\tau}}\}.\end{split}$ (7)
@@
Note that for any trajectory $X^{\mu}(\tau)$, there exist infinitely many spacetime metrics $g_{\alpha\beta}$ with respect to which this trajectory can be viewed as a geodesic. Spacetimes that satisfy this condition are defined to belong to the same class of spacetime, denoted by $g_{\alpha\beta}]$, whereas spacetimes in which this trajectory cannot be regarded as a geodesic do not belong to this class. This implies that each trajectory appearing in the path integral (4) corresponds to a distinct class of spacetimes $g_{\alpha\beta}]$. For a specific spacetime, it belongs to, and only to, one of the classes $g_{\alpha\beta}]$. As each trajectory is only counted once in the path integral (4), in each class of spacetime $g_{\alpha\beta}]$, only one metric contributes to the path integral. However, in Equation (7), the integral $\int Dg_{\mu\nu}$ contains all metrics for every class of spacetime $g_{\alpha\beta}]$. We need to eliminate the redundant metrics.
@@
$\mathcal{G}^{(m)}(g_{\alpha\beta})=0,\ \ \ \ \ \ \ \ \ m=1,\ 2,\ \ldots,\ M.$ (8)
@@
Here, $M$ represents the total number of gauge conditions. Similar to the scenario with Yang–Mills gauge fields, the gauge conditions have a certain degree of arbitrariness. As long as the solution to Equation (8) can uniquely determine a metric from each class of spacetime $g_{\alpha\beta}]$, the gauge conditions are considered valid. For every trajectory in the path integral (4), there exists a single metric that satisfies both Equations (5) and (8).
@@
According to the Faddeev–Popov method [31, 32], to eliminate the redundant metrics, Equation (7) should be modified as
@@
$\langle X^{\mu}(\tau_{f})|X^{\mu}(\tau_{i})\rangle=N\int Dg_{\mu\nu}\ \mathcal{J}\ \Delta(\mathcal{G}^{(m)})\prod_{m=1}^{M}\delta(\mathcal{G}^{(m)})\ \mathrm{exp}\{im_{0}\int_{\tau_{i}}^{\tau_{f}}d\tau\sqrt{g_{\alpha\beta}\frac{d\mathcal{F}^{\alpha}}{d\tau}\frac{d\mathcal{F}^{\beta}}{d\tau}}\}.$ (9)
@@
Here, $\Delta(\mathcal{G}^{(m)})$ and $\delta(x)$ represent the Faddeev–Popov determinant and the $\delta$-function, respectively. The Jacobian determinant $\mathcal{J}$ is determined by Equation (3), while the Faddeev–Popov determinant depends on the gauge conditions (8). The conclusions of this work are independent of the explicit form of $\mathcal{J}$, $\Delta(\mathcal{G}^{(m)})$ and $\mathcal{F}^{\mu}$. Thus, we will avoid solving them. This is a challenging issue that needs careful study in the future.
@@
Equation (9) represents a new kind of representation of quantum mechanics. We name it the spacetime representation. Equation (4) is equivalent to Equation (9). Thus, the integration over different trajectories can be transformed into integration over various spacetime metrics. Equation (9) indicates that the quantum characteristics of the particle can be attributed to the uncertainty of the spacetime structure for the microscopic particle. In the conventional interpretation of path integral quantum mechanics, the spacetime structure is deterministic, and only the classical trajectory corresponds to the Hamilton–Jacobi equation. All other trajectories are not solutions of any dynamical equations. Thus, the classical trajectory possesses a special status. However, we argued that for the free particle, all trajectories can be seen as the geodesics in different spacetimes. Thus, different trajectories are solutions of the geodesic equation in different spacetimes. This perspective offers a more natural understanding of these trajectories. In Table 1, we outlined the key distinctions between the conventional path integral and the spacetime representation for a single relativistic free particle. When $m_{0}\rightarrow\infty$, the microscopic particle becomes a macroscopic object. The coherence disappears and the dynamics of the particle can be described by classical physics. Consequently, the spacetime structure becomes deterministic for macroscopic objects.
  Item Conventional path integral Spacetime representation
  Partition function Equation (4) Equation (9)
  Integral variable Particle’s trajectory Spacetime metric
  Spacetime structure Spacetime structure is deterministic (Minkowski spacetime). Spacetime structure is uncertain for the microscopic particle.
@@
  Particle’s trajectories Only the classical trajectory is the solution of the Hamilton-Jacobi equation. All other trajectories in the path integral (4) do not correspond to any dynamical equation. All trajectories in the path integral (4) are viewed as geodesics in different spacetimes.